OpenStudy (anonymous):

Find the solution to the system of equations by substitution 12x-y=1 20x+1/3y=1

OpenStudy (anonymous):

\[y=12x-1.........and....... y=-60x+3 .....after you multiply... by3\] now solve 72x=4 x=4/72 =1/18 put in first equation y =-1/3